### v1

# Quasi-random hypergraphs arising from multivariate polynomials and their applications to Diophantine sets

Abstract.
In this paper, we provide new explicit constructions of families of quasi-random hypergraphs using algebraic properties of multivariate polynomials over finite fields. Our construction not only offers a systematic method for constructing quasi-random hypergraphs but also provides a unified framework for studying various hypergraphs arising from multivariate polynomials over finite fields, including Paley sum hypergraphs, and hypergraphs derived from Diophantine tuples and their generalizations. In particular, we refine an estimate related to even partial octahedrons in Paley sum hypergraphs, improving upon a result by Chung and Graham. Inspired by our quasi-random constructions, we establish an asymptotic formula for the number of $k$-Diophantine $m$-tuples over finite fields, addressing a question of Hammonds et al.; we also study some related questions for $f$-Diophantine sets, extending and improving several recent works.
Key words and phrases:
quasi-random hypergraph, polynomial, Diophantine tuple
2020 Mathematics Subject Classification:
Primary: 11T06, 05C80. Secondary: 05C65, 11D72, 05C25, 11B30
1. Introduction
Throughout the paper, let $k\geq 2$ be a fixed integer. Let $q$ be an odd prime power, ${\mathbb{F}}_{q}$ the finite field with $q$ elements, and ${\mathbb{F}}_{q}^{*}={\mathbb{F}}_{q}\setminus\{0\}$. Let $\overline{{\mathbb{F}}_{q}}$ be the algebraic closure of ${\mathbb{F}}_{q}$.
The concept of quasi-randomness for graphs was introduced in [10], and independently in [32, 33]. In particular, Chung, Graham, and Wilson [10] showed that several seemingly unrelated properties of quasi-random graphs are, in fact, equivalent. Building on this, Chung and Graham [7, 9] extended the concept of quasi-randomness to uniform hypergraphs, maintaining a similar perspective as for quasi-random graphs. In particular, they showed that when the edge density of a family of hypergraphs is $1/2$, several analogous statements of the properties of quasi-random graphs are asymptotically equivalent. Since the work of Chung and Graham, various approaches to generalizing the concept of quasi-randomness for hypergraphs have been explored in the literature [1, 7, 11, 16, 21, 22, 25, 26, 35], addressing different directions and applications. For the purpose of our paper, we adopt the notion of quasi-randomness defined by Chung and Graham [9] as their definitions include subgraph counting for general hypergraphs.
We first introduce some necessary notations to define quasi-random hypergraphs. Let $G$ and $H$ be $k$-uniform hypergraphs. We use $N^{*}_{G}(H)$ to denote the number of labeled occurrences of $H$ as an induced subgraph of $G$. For a family $\mathcal{F}$ of hypergraphs, we denote $N_{G}^{*}(\mathcal{F})=\sum_{H\in\mathcal{F}}N_{G}^{*}(H)$. Next, we define a special type of uniform hypergraph that plays a key role in defining quasi-random hypergraphs. A $k$-uniform hypergraph is called a $k$-octahedron, denoted by
$\mathcal{O}_{k}=\mathcal{O}_{k}(u_{1}(0),u_{1}(1),u_{2}(0),u_{2}(1),\ldots,u_{k}(0),u_{k}(1)),$
where the vertex set consists of $u_{i}(\varepsilon_{i})$, $1\leq i\leq k$, $\varepsilon_{i}\in\{0,1\}$. The edge set is defined as the set of all $k$-sets of the form $\{u_{1}(\varepsilon_{1}),u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k})\mid\varepsilon_{i}\in\{0,1\},1\leq i\leq k\}$. In other words, a $k$-octahedron is a complete $k$-uniform $k$-partite hypergraph where each vertex class has two vertices. Further, we say a $k$-uniform hypergraph is an even partial octahedron (EPO) if it has the same vertex set as $\mathcal{O}_{k}$, and the intersection of its edge set and the edge set of $\mathcal{O}_{k}$ has even size. The set of all EPOs defined on $\mathcal{O}_{k}$ is denoted by $\mathcal{O}_{k}^{e}$.
Following Chung and Graham [9], we introduce the following properties of $k$-uniform hypergraphs.
Definition 1.1.
[9, Section 3] Let $\mathcal{G}$ be a family of $k$-uniform hypergraphs with an unbounded number of vertices; we say $\mathcal{G}$ has a property below if all $G\in\mathcal{G}$ satisfy the property.
  1. $Q_{1}(s)$ :
    For all $k$-uniform hypergraphs $G^{\prime}$ with $s$ vertices,
    $N^{*}_{G}(G^{\prime})=(1+o(1))\frac{n^{s}}{2^{\binom{s}{k}}},\quad\text{ as }|V(G)|=n\to\infty.$
  2. $Q_{3}$ :
    The number of induced copies of EPOs in $G$ is
    $N^{*}_{G}(\mathcal{O}_{k}^{e})\leq(1+o(1))\frac{n^{2k}}{2},\quad\text{ as }|V(G)|=n\to\infty.$
The main result in [9] states that if $s\geq 2k$, then $Q_{1}(s)$ and $Q_{3}$ are equivalent  [^1] On the other hand, as remarked in [9, Section 11], $Q_{1}(s)$ is not equivalent to $Q_{3}$ for $s\leq k+1$. . Following [9], we say $\mathcal{G}$ is a family of quasi-random ($k$-uniform) hypergraphs if $\mathcal{G}$ has property $Q_{3}$. For the purpose of this paper, we also introduce the following two additional equivalent properties, $Q_{1}$ and $Q_{3}^{\prime}$, that can be used to define quasi-random hypergraphs.
  • $Q_{1}$ :
    $Q_{1}(s)$ holds for all positive integers $s$.
  • $Q_{3}^{\prime}$ :
    The number of induced copies of EPOs in $G$ is
    $N^{*}_{G}(\mathcal{O}_{k}^{e})=(1+o(1))\frac{n^{2k}}{2},\quad\text{ as }|V(G)|=n\to\infty.$
It is easy to show that for each positive integer $s$, $Q_{1}(s+1)$ implies $Q_{1}(s)$ [9, page 108]. It follows that $Q_{1}$ and $Q_{3}$ are equivalent, since, as mentioned earlier, $Q_{3}$ is equivalent to $Q_{1}(s)$ for each $s\geq 2k$. The equivalence of $Q_{3}$ and $Q_{3}^{\prime}$ can be found in [9, Fact 2].
Numerous explicit constructions of quasi-random graphs have been extensively studied. Most of these constructions belong to the classes of either Cayley graphs or Cayley sum graphs. Classical examples include cyclotomic graphs and cyclotomic sum graphs, which are well-established as Cayley or Cayley sum quasi-random graphs [5, 8, 10, 20]. Recently, the authors [18] introduced a novel class of quasi-random graphs that are neither the Cayley or Cayley sum graphs using polynomials over finite fields. For further explicit constructions of quasi-random graphs, we refer the readers to [4, 32, 34]. We also refer the readers to an excellent survey on pseudo-random graphs by Krivelevich and Sudakov [23].
Despite the importance of the quasi-randomness of hypergraphs in combinatorics, explicit families of quasi-random hypergraphs remain largely unexplored. One of the classical examples of quasi-random hypergraphs is the family of Paley (sum) hypergraphs; see [9, Section 10]. For an odd prime power $q$, the Paley (sum) $k$-graph $P_{k}(q)$ is the $k$-uniform hypergraph whose vertices are the elements of ${\mathbb{F}}_{q}$, such that a $k$-set $\{u_{1},u_{2},\ldots,u_{k}\}$ forms an edge if and only if $u_{1}+u_{2}+\cdots+u_{k}$ is a square in ${\mathbb{F}}_{q}$. Other than Paley $k$-graphs and even intersection $k$-graphs in [9, Section 10], we refer to [2] for other examples of such families.
The main goal of this paper is to provide a new systematic approach to constructing explicit families of quasi-random hypergraphs using multivariate polynomials over finite fields. By using the quasi-randomness of these new constructions, we also resolve a question posed in [15, Section 4] related to Diophantine equations. Specifically, we derive an asymptotic formula for the number of $k$-Diophantine $m$-tuples over finite fields; see Subsection 1.2 for a more precise statement.
1.1. Main results
Let $q$ be an odd prime power and let $f\in\mathbb{F}_{q}[x_{1},x_{2},\ldots,x_{k}]$ be a symmetric polynomial. We define the $k$-uniform hypergraph $Y_{f,q}$ on ${\mathbb{F}}_{q}$, such that a $k$-element subset $\{a_{1},a_{2}\ldots,a_{k}\}$ of ${\mathbb{F}}_{q}$ forms an edge if and only if $f(a_{1},a_{2},\ldots,a_{k})$ is a square in ${\mathbb{F}}_{q}$.
Definition 1.2.
Let $f\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$ be a symmetric polynomial with degree $d$. We say that $f$ is admissible if it satisfies the following two conditions:
  • $f(x_{1},x_{2},\ldots,x_{k})$ is not a constant multiple of the square of a polynomial in ${\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$.
  • $f$ satisfies the primitive condition: when we write
    $f(x_{1},x_{2},\ldots,x_{k})=\sum_{j=0}^{d}H_{j}(x_{2},x_{3},\ldots,x_{k})x_{1}^{j},$
    the polynomials $H_{0},H_{1},H_{2},\ldots,H_{d}$ do not share a common zero over $\overline{{\mathbb{F}}_{q}}$.
We now present our main result.
Theorem 1.3.
Let $k\geq 2$ and let $d\geq 1$. Let $\mathcal{Y}_{d}$ be the family consisting of all hypergraphs $Y_{f,q}$, where $q$ is an odd prime power and $f\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$ is an admissible polynomial with degree $d$. Then
$N^{*}_{Y_{f,q}}(\mathcal{O}_{k}^{e})=\frac{q^{2k}}{2}+O(q^{2k-1})$
holds uniformly for all $Y_{f,q}\in\mathcal{Y}_{d}$. In particular, $\mathcal{Y}_{d}$ is a family of quasi-random $k$-uniform hypergraphs.
As a direct consequence of Theorem 1.3, we have the following corollary.
Corollary 1.4.
The number of induced copies of EPOs in $P_{k}(q)$ is $q^{2k}/2+O(q^{2k-1}).$ In particular, the family of Paley $k$-graphs is quasi-random.
In particular, Corollary 1.4 improves the error term in the previous result $N_{P_{k}(q)}^{*}(\mathcal{O}_{k}^{e})=q^{2k}/2+O(q^{2k-1/2})$ proved by Chung and Graham [9, Section 10].
In Section 2, we provide further motivation for the two conditions used to define admissible polynomials. These two conditions are fundamental. Indeed, if the first condition is not satisfied, then the graph $Y_{f,q}$ would be far from being quasi-random; see Remark 2.2. While the second condition seems restrictive, it is not hard to show that if $d$ and $k$ are fixed, then almost all symmetric polynomials in ${\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$ with degree $d$ are admissible, as $q\to\infty$; see Remark 2.4. Thus, informally speaking, almost all $Y_{f,q}$ are quasi-random. In addition, we compare our second condition with the notion “primitive kernel” for $2$-variable polynomials introduced by Gyarmati and Sárközy [14]; see Remark 2.3.
1.2. Applications to Diophantine tuples over finite fields
Finally, we discuss some applications of Theorem 1.3 to Diophantine tuples. A set of distinct $m$ positive integers is called a Diophantine $m$-tuple if the product of any two distinct elements in the set is one less than a perfect square. There are many well-studied variants of Diophantine tuples, and we refer to a recent book by Dujella [12] for a comprehensive discussion. For our purposes, we consider finite field analogues of Diophantine tuples.
Let $k\geq 2$, $q$ be an odd prime power, and $f\in{\mathbb{F}}_{q}[x_{1},\ldots,x_{k}]$ be a symmetric polynomial. Following [3, 37], we say $A\subset{\mathbb{F}}_{q}$ is an $f$-Diophantine set over ${\mathbb{F}}_{q}$ if $f(a_{1},a_{2},\ldots,a_{k})$ is a square in ${\mathbb{F}}_{q}$ whenever $a_{1},a_{2},\ldots,a_{k}$ are distinct elements in $A$. In particular, when $k=2$ and $f(x_{1},x_{2})=x_{1}x_{2}+1$, an $f$-Diophantine set over ${\mathbb{F}}_{q}$ is exactly a Diophantine tuple over ${\mathbb{F}}_{q}$, which is of special interest [13, 17, 18, 19, 28, 30, 36]. More generally, when $k\geq 2$ and $f(x_{1},x_{2},\ldots,x_{k})=x_{1}x_{2}\cdots x_{k}+1$, following Hammonds et al. [15], an $f$-Diophantine set with size $m$ over ${\mathbb{F}}_{q}$ is said to be a $k$-Diophantine $m$-tuple over ${\mathbb{F}}_{q}$. Theorem 1.3 implies the following corollary immediately.
Corollary 1.5.
The family of $k$-graphs induced by $k$-Diophantine tuple is quasi-random. More precisely, if $k\geq 2$, then the family of $k$-graphs $\mathcal{D}_{k}=\{Y_{f,q}:q\text{ is an odd prime power}\}$ is quasi-random, where $f(x_{1},x_{2},\ldots,x_{k})=x_{1}x_{2}\cdots x_{k}+1$.
Hammonds et al. [15, Section 4] asked the following question: Can we count the number of $k$-Diophantine $m$-tuples in ${\mathbb{F}}_{q}$ for a given odd prime power $q$? While an explicit formula is believed to be out of reach in general, an asymptotic formula for the case $k=m$ was answered in [15, Theorem 1.7], and the general case $m>k$ was left as an open problem.
We resolve their question in the following corollary. Corollary 1.5 implies that the family $\mathcal{D}_{k}$ has the property $Q_{1}$, so the corollary follows immediately.
Corollary 1.6.
Let $k$ and $m$ be fixed positive integers with $2\leq k\leq m$. Let $q$ be an odd prime power. Then as $q\to\infty$, the number of $k$-Diophantine $m$-tuples in ${\mathbb{F}}_{q}$ is $q^{m}/\big{(}m!\cdot 2^{\binom{m}{k}}\big{)}+o(q^{m}).$
Note that when $m=k$, the above corollary recovers [15, Theorem 1.7]. We remark that the case $k=2$ in Corollary 1.6 has been proved by Dujella and Kazalicki [13, Section 6]; however, as remarked in [15, Section 4], it appears that the techniques in [13, 15] do not extend to $k$-Diophantine tuples with $k\geq 3$ and $m>k$.
Our proof of Theorem 1.3 (and thus Corollary 1.6) only relies on the Weil’s bound on complete character sums (with single variable polynomial argument, see Lemma 4.2) and the main result in Chung and Graham [9]. However, since we are dealing with multivariate polynomials and we already have an asymptotic formula, naturally one expects that there is a power saving in the error term by appealing to more sophisticated higher dimensional versions of Weil’s bound. Indeed, using a consequence of the Lang-Weil bound for geometrically irreducible varieties [24], in Section 5, we obtain the following effective version of Corollary 1.6 that holds for all admissible polynomials $f$.
Theorem 1.7.
Let $k\geq 2$ be a fixed integer. Then uniformly for all odd prime powers $q$, all admissible polynomial $f\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$ with degree $d\geq 1$ and all integers $m\geq k$, the number of $f$-Diophantine sets of size $m$ is
$\frac{q^{m}}{m!\cdot 2^{\binom{m}{k}}}+O\big{(}d^{2\binom{m}{k}}q^{m-1/2}+(2d)^{13\binom{m}{k}/3}q^{m-1}\big{)}.$
We remark that Shparlinski [30] showed that the error term in Theorem 1.7 can be improved to $O_{m}(q^{m-1})$ when $k=2$ and $f(x_{1},x_{2})=x_{1}x_{2}+r$ for $r\in{\mathbb{F}}_{q}^{*}$. However, his method seems to rely on the structure of the polynomial $x_{1}x_{2}+r$ and does not seem to extend to general polynomials.
It is well-known that a random $k$-graph $G_{k}(n,1/2)$ has clique number of the order $(\log n)^{1/(k-1)}$ almost surely. In view of Theorem 1.3, whenever $f\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$ is admissible with degree $d$, one expects that $\omega(Y_{f,q})$, the clique number of $Y_{f,q}$, equivalently, the largest size of $f$-Diophantine sets over ${\mathbb{F}}_{q}$, to be at least $c_{d,k}(\log q)^{1/(k-1)}$, where $c_{d,k}$ is a positive constant depending only on $d$ and $k$. When $k=2$, this heuristic has been confirmed in [18, Theorem 1.4 and Theorem 1.6] by introducing a quantitative way to measure the quasi-randomness for graphs. However, studying cliques in hypergraphs is usually much harder compared to that in graphs, and it seems challenging to extend the method in [18] to hypergraphs.
Theorem 1.7 readily provides the following lower bound on the clique number of $Y_{f,q}$, which, while not necessarily optimal, represents an improvement on some previous results.
Corollary 1.8.
If $k\geq 2$, $q$ is an odd prime power, and $f\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$ is an admissible polynomial with degree $d\geq 1$, then $\omega(Y_{f,q})\geq c_{k}(\log_{\widetilde{d}}q)^{1/k}$, where $\widetilde{d}=\max(2,d)$, and $c_{k}$ is a positive constant only depending on $k$.
We remark that if $f$ is a sparse polynomial, a better lower bound on $\omega(Y_{f,q})$ was proved in [37] based on some explicit algebraic constructions. More precisely, under some additional structural assumptions on $f$, it is proved in [37, Theorem 1.1] that $\omega(Y_{f,q})\geq(1-o(1))\frac{1}{d}(\log_{4}q)^{1/\ell(f)}$, where $\ell(f)$ is the number of non-constant monomials in the monomial expansion of $f$ and $d$ is the degree of $f$. In particular, $\omega(Y_{f,q})\geq(\frac{1}{k}-o(1))\log_{4}q$ when $f=x_{1}x_{2}\ldots x_{k}+1$. However, note that the parameter $\ell(f)$ is much larger compared to $k$ for a generic polynomial $f$, and thus the above corollary refines this result for almost all admissible polynomials.
Notations. We follow the standard asymptotic notations. We use the standard big $O$ notation, meaning the implicit constant is absolute. $d$ always stands for the degree of a polynomial, and $O_{d}$ means the implicit constant only depends on the degree $d$ (but not on the polynomial itself).
Structure of the paper. In Section 2, we provide more backgrounds. In Section 3, we prove a technical proposition that is necessary for proving our main result. In Section 4, we prove our main result, Theorem 1.3. Finally, we prove Theorem 1.7 in Section 5.
2. More backgrounds and motivations
In this section, we give more backgrounds and motivations for defining admissible polynomials (Definition 1.2).
We first recall the following fundamental result, which is a special case of the Schwartz–Zippel lemma [29, Corollary 1] applied to finite fields. We will apply the lemma repeatedly in the paper.
Lemma 2.1 (Schwartz–Zippel lemma).
Let $d\geq 1$ and let $g\in\overline{{\mathbb{F}}_{q}}[x_{1},x_{2},\ldots,x_{m}]$ be a polynomial with degree $d$. Then the number of $(u_{1},u_{2},\ldots,u_{m})\in{\mathbb{F}}_{q}^{m}$ such that $g(u_{1},u_{2},\ldots,u_{m})=0$ is at most $dq^{m-1}$.
Remark 2.2.
Assume that $f(x_{1},x_{2},\ldots,x_{k})$ is a constant multiple of the square of a polynomial. In this case, $Y_{f,q}$ is far from being quasi-random. Say $f=cg^{2}$, where $c\in{\mathbb{F}}_{q}$ and $g\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$. If $c$ is a square in ${\mathbb{F}}_{q}$, then by definition, $Y_{f,q}$ is complete. On the other hand, if $c$ is a non-square in ${\mathbb{F}}_{q}$, then $k$ distinct vertices $\{v_{1},v_{2},\ldots,v_{k}\}$ form an edge in $Y_{f,q}$ if and only if $f(v_{1},v_{2},\ldots,v_{k})=0$. Thus, by the Schwartz–Zippel lemma, the number of edges of $Y_{f,q}$ is $O_{d}(q^{k-1})$ and thus $Y_{f,q}$ is sparse.
Next, we give two remarks related to the primitive condition in the definition of admissible polynomials.
Remark 2.3.
A polynomial $f(x,y)\in{\mathbb{F}}_{q}[x,y]$ can be expressed as
$f(x,y)=\sum_{i=0}^{m}G_{i}(y)x^{i}=\sum_{j=0}^{n}H_{j}(y)y^{j}.$
Following [14], we say $f$ is primitive in $x$ if $\gcd(G_{0}(y),\ldots,G_{m}(y))=1$, and primitive in $y$ if $\gcd(H_{0}(y),\ldots,H_{m}(y))=1$. In [14], Gyarmati and Sárközy proved that any two-variable polynomial $f(x,y)\in\mathbb{F}_{q}$ has a factorization, which is unique up to constant factors, such that $f(x,y)=F(x)G(y)H(x,y)$, where $H(x,y)$ is the primitive kernel of $f(x,y)$ in the sense that $H(x,y)$ is primitive in both $x$ and $y$.
Next, we show that our primitive condition in Definition 1.2 extends their primitivity notion to polynomials with $k\geq 2$ variables. Recall we say that $f\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$ satisfies the primitive condition provided that when we write $f(x_{1},x_{2},\ldots,x_{k})=\sum_{j=0}^{d}H_{j}(x_{2},x_{3},\ldots,x_{k})x_{1}^{j},$ the polynomials $H_{0},H_{1},H_{2},\ldots,H_{d}$ do not share a common zero over $\overline{{\mathbb{F}}_{q}}$. Since we work on symmetric polynomials, the same condition holds if we expand $f$ based on each $x_{i}$. Note that if $k=2$, this primitive condition is equivalent to $\gcd(H_{0},H_{1},\ldots,H_{d})=1$, which is consistent with the above notion of primitivity of $2$-variable polynomials.
However, in general, when $k\geq 3$, the condition $\gcd(H_{0},H_{1},\ldots,H_{d})=1$ is not equivalent to the condition that $H_{0},H_{1},H_{2},\ldots,H_{d}$ do not share a common zero over $\overline{{\mathbb{F}}_{q}}$. For example, if $f(x_{1},x_{2},x_{3})=x_{1}x_{2}+x_{2}x_{3}+x_{3}x_{1}=H_{1}(x_{2},x_{3})x_{1}+H_{2}(x_{2},x_{3})$ with $H_{1}(x_{2},x_{3})=x_{2}+x_{3}$ and $H_{2}(x_{2},x_{3})=x_{2}x_{3}$, then $\gcd(H_{1},H_{2})=1$, however, $H_{1}$ and $H_{2}$ share the common root $x_{2}=x_{3}=0$. The same example also shows that the factorization result mentioned above does not extend in general to the $3$ or more variables case: the polynomial $f(x_{1},x_{2},x_{3})$ does not satisfy the primitive condition, however, it is irreducible over $\mathbb{F}_{q}$.
Remark 2.4.
One can prove that almost all multivariable symmetric polynomials of any degree are admissible over $\mathbb{F}_{q}$. We briefly illustrate this for the case of $3$-variable polynomials of degree $2$. We can write any symmetric $f(x,y,z)$ of degree $2$ as
$f(x,y,z)=Ax^{2}+Ay^{2}+Az^{2}+Bxy+Byz+Bxz+Cx+Cy+Cz+D.$ (2.1)
We let $F_{0}=F_{0}(y,z)=D+Cy+Cz+Byz+Ay^{2}+Az^{2}$ and $F_{1}=F_{1}(y,z)=By+Bz+C$ so that
$f(x,y,z)=Ax^{2}+F_{1}(y,z)x+F_{0}(y,z).$
When $A\neq 0$, we can guarantee that $f(x,y,z)$ satisfies the primitive condition. Now we assume $A=0$. Then we have $f(x,y,z)=F_{0}(y,z)+F_{1}(y,z)x$, and we just need to determine the cases when $F_{0}$ and $F_{1}$ share common zeros or not to determine the primitivity of $f(x,y,z)$. That is to say, whether their resultant is zero or not. The resultant of $F_{0}$ and $F_{1}$ in terms of the variable $y$ is
$\textrm{Res}(F_{0},F_{1},y)=\det\begin{bmatrix}Bz+C&B\\ D+Cz&C+Bz\end{bmatrix}=B^{2}z^{2}+BCz+C^{2}-BD.$
Hence $\textrm{Res}(F_{0},F_{1},y)$ has no roots in $\overline{{\mathbb{F}}_{q}}$ if and only if $B=0$ and $C^{2}-BD\neq 0$, thus if and only if $B=0$, and $C\neq 0$ (If $C=0,D=0$, then the resultant is zero when $y=0$ for any $B$. This is the case of $xy+yz+zx$ in Remark 2.3.). Hence, the probability of a randomly chosen polynomial 2.1 satisfying the primitive condition is
$\frac{(q-1)q^{3}+(q-1)q}{q^{4}}\to 1\quad\text{as $\quad q\to\infty$.}$
On the other hand, it is easy to show that the probability of a randomly chosen polynomial 2.1 being a constant multiple of the square of a polynomial in ${\mathbb{F}}_{q}[x,y,z]$ tends to 00 as $q\to\infty$.
The same argument shows that if $d$ and $k$ are fixed, then almost all symmetric polynomials in ${\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$ with degree $d$ are admissible, as $q\to\infty$. Indeed, to see that almost all such polynomials satisfy the primitive condition, it suffices to consider the coefficient for $x_{1}^{d}$.
3. A key proposition
The main purpose of this section is to prove Proposition 3.1, which will play a key role in Section 4. Our proof adapts the approach of Gyarmati and Sárközy [14, Section 6], extending their results from the $2$-variable case to the $k$-variable case for any $k\geq 2$.
Proposition 3.1.
Let $q$ be odd prime power and $k\geq 2$. Assume that $f(x_{1},x_{2},\ldots,x_{k})\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$ is admissible with degree $d$. Denote by $X$ the set of tuples $(u_{2},u_{3},\ldots,u_{k})\in{\mathbb{F}}_{q}^{k-1}$ with $f(x_{1},u_{2},u_{3},\ldots,u_{k})=ch^{2}(x_{1})$ for some $c\in{\mathbb{F}}_{q}$ and monic $h(x_{1})\in{\mathbb{F}}_{q}[x_{1}]$. Then we have $|X|\leq(d^{2}+d)q^{k-2}.$
Write
$f(x_{1},x_{2},\ldots,x_{k})=p_{n}(x_{2},\ldots,x_{k})x_{1}^{n}+\cdots+p_{1}(x_{2},\ldots,x_{k})x_{1}+p_{0}(x_{2},\ldots,x_{k}),$ (3.1)
where $p_{n}$ is not identically zero. Note that $n\geq 1$; otherwise, $f(x_{1},x_{2},\ldots,x_{k})$ is a constant polynomial, violating the assumption. We define
$Y$ =$=\{(u_{2},\ldots,u_{k})\in X\mid p_{n}(u_{2},\ldots,u_{k})=0\},\quad\text{and}$
$Z$ =$=\{(u_{2},\ldots,u_{k})\in X\mid p_{n}(u_{2},\ldots,u_{k})\neq 0\}.$
Then $|X|=|Y|+|Z|$. By the Schwartz–Zippel lemma, we have $|Y|\leq(d-n)q^{k-2}\leq dq^{k-2}$. We now proceed to estimate $|Z|$. The following lemma completes the proof of Proposition 3.1.
Lemma 3.2.
We have $|Z|\leq ndq^{k-2}.$ In particular, $|Z|\leq d^{2}q^{k-2}$.
Proof.
If $n$ is odd, then clearly $Z=\emptyset$ and we are done. Next, assume that $n$ is even.
If $(u_{2,1},u_{3,1},\ldots,u_{k,1}),\ldots,(u_{2,T},u_{3,T},\ldots,u_{k,T})$ are all the distinct elements of $Z$, then for each $i=1,\ldots,T$, the polynomials $f(x_{1},u_{2,i},u_{3,i},\ldots,u_{k,i})$ can be written in the form of
$f(x_{1},u_{2,i},u_{3,i},\ldots,u_{k,i})=C_{i}h^{2}_{i}(x_{1}),\quad\text{with $C_{i}\in{\mathbb{F}}_{q}$ and monic $h_{i}(x_{1})\in{\mathbb{F}}_{q}[x_{1}].$}$
Note that $C_{i}=p_{n}(u_{2,i},u_{3,i},\ldots,u_{k,i})$ and hence
$f(x_{1},u_{2,i},u_{3,i},\ldots,u_{k,i})=p_{n}(u_{2,i},u_{3,i},\ldots,u_{k,i})h^{2}_{i}(x_{1}).$ (3.2)
We denote
$h_{i}(x_{1})=a_{r}(i)x_{1}^{r}+a_{r-1}(i)x_{1}^{r-1}+\cdots+a_{1}(i)x_{1}+a_{0}(i),$ (3.3)
where $a_{r}(i)=1$, $a_{r-j}(i)\in{\mathbb{F}}_{q}$, and $r=\frac{n}{2}$.
Claim 3.3.
For each $0\leq j\leq r$, there exists a polynomial $q_{r-j}(x_{2},x_{3},\ldots,x_{k})\in{\mathbb{F}}_{q}[x_{2},x_{3},\ldots,x_{k}]$ with degree at most $jd$, such that
$a_{r-j}(i)(p_{n}(u_{2,i},u_{3,i},\ldots,u_{k,i}))^{j}=q_{r-j}(u_{2,i},u_{3,i},\ldots,u_{k,i})$ (3.4)
holds for $1\leq i\leq T$.
Proof of claim.
We prove the claim by induction. The case $j=0$ is obvious since we can simply set $q_{r}\equiv 1$. Let $1\leq j\leq r$ and assume that the claim holds for $10,1,\ldots j-$ 1. The coefficient of $x_{1}^{n-j}$ in $p_{n}(u_{2,i},u_{3,i},\ldots,u_{k,i})^{j}h^{2}_{i}(x_{1})$ is
$(p_{n}(u_{2,i},\ldots,u_{k,i}))^{j}\left(2a_{r-j}(i)+\sum_{\begin{subarray}{c}0\leq t_{0},\ldots t_{j-1}\\ t_{0}+\cdots+t_{j-1}=2\\ t_{1}+2t_{2}+\cdots+(j-1)t_{j-1}=j\end{subarray}}\frac{2}{t_{0}!\ldots t_{j-1}!}a_{r-1}(i)^{t_{1}}\cdots a_{r-(j-1)}(i)^{t_{j-1}}\right)$
$=2a_{r-j}(i)(p_{n}(u_{2,i},\ldots,u_{k,i}))^{j}$
$+\sum_{\begin{subarray}{c}0\leq t_{0},\ldots t_{j-1}\\ t_{0}+\cdots+t_{j-1}=2\\ t_{1}+2t_{2}+\cdots+(j-1)t_{j-1}=j\end{subarray}}\frac{2}{t_{0}!\ldots t_{j-1}!}(a_{r-1}(i)p_{n}(u_{2,i},\ldots,u_{k,i}))^{t_{1}}\cdots(a_{r-(j-1)}(i)p_{n}(u_{2,i},\ldots,u_{k,i})^{j-1})^{t_{j-1}},$
which, by the inductive hypothesis, is of the form
$2a_{r-j}(i)(p_{n}(u_{2,i},u_{3,i},\ldots,u_{k,i}))^{j}+s_{j}(u_{2,i},u_{3,i},\ldots,u_{k,i}),$
where $s_{j}(x_{2},x_{3},\ldots,x_{k})\in{\mathbb{F}}_{q}[x_{2},x_{3},\ldots,x_{k}]$ is a polynomial (independent of $i$) of degree $\leq jd$. Multiplying $(p_{n}(u_{2,i},u_{3,i},\ldots,u_{k,i}))^{j-1}$ on both sides of equation (3.2) and comparing the coefficient of $x_{1}^{n-j}$ , we have the equality
$p_{n-j}(u_{2,i},\ldots,u_{k,i})(p_{n}(u_{2,i},\ldots,u_{k,i}))^{j-1}=2a_{r-j}(i)(p_{n}(u_{2,i},\ldots,u_{k,i}))^{j}+s_{j}(u_{2,i},\ldots,u_{k,i}),$
which implies
$a_{r-j}(i)(p_{n}(u_{2,i},\ldots,u_{k,i}))^{j}=\frac{1}{2}(p_{n-j}(u_{2,i},\ldots,u_{k,i})(p_{n}(u_{2,i},\ldots,u_{k,i}))^{j-1}-s_{j}(u_{2,i},\ldots,u_{k,i})).$
Thus, we can set
$q_{r-j}(x_{2},\ldots,x_{k})=\frac{1}{2}\big{(}p_{n-j}(x_{2},\ldots,x_{k})(p_{n}(x_{2},\ldots,x_{k}))^{j-1}-s(x_{2},\ldots,x_{k})\big{)},$
with degree at most $dj$ such that equation (3.4) holds, as required. ∎
We define
$h$ $(x_{1},x_{2},\ldots,x_{k})$
$=(p_{n}(x_{2},\ldots,x_{k}))^{r}x_{1}^{r}+q_{r-1}(x_{2},\ldots,x_{k})(p_{n}(x_{2},\ldots,x_{k}))^{r-1}x_{1}^{r-1}+\cdots+q_{0}(x_{2},\ldots,x_{k}),$
where the polynomials $q_{r-j}$’s are from 3.3 and the degree of $q_{r-j}$ is at most $jd$. Then by equations (3.3) and (3.4), we have
$(p_{n}(u_{2,i},\ldots,u_{k,i}))^{r}h_{i}(x_{1})=h(x_{1},u_{2,i},\ldots,u_{k,i})$
for $1\leq i\leq T$. It then follows from equation (3.2) that
$(p_{n}(u_{2,i},\ldots,u_{k,i}))^{n-1}f(x_{1},u_{2,i},\ldots,u_{k,i})=h^{2}(x_{1},u_{2,i},\ldots,u_{k,i})$ (3.5)
holds for each $1\leq i\leq T$.
Let $1\leq\ell\leq n$. From the definition of $h(x_{1},x_{2},\ldots,x_{k})$ and (3.5), we can deduce that the coefficient of $x_{1}^{n-\ell}$ of $(p_{n}(u_{2,i},\ldots,u_{k,i}))^{n-1}f(x_{1},u_{2,i},\ldots,u_{k,i})$ is
$r_{n-\ell}(u_{2,i},\ldots,u_{k,i})(p_{n}(u_{2,i},\ldots,u_{k,i}))^{n-\ell},$
where $r_{n-\ell}(x_{2},\ldots,x_{k})\in{\mathbb{F}}_{q}[x_{2},\ldots,x_{k}]$ is a polynomial of degree at most $\ell d$.
Multiplying both sides of equation (3.1) by $(p_{n}(x_{2},\ldots,x_{k}))^{n-1}$ and comparing the coefficient of $x_{1}^{n-\ell}$, for each $1\leq i\leq T$, we deduce from equation (3.5) that
$p_{n-\ell}(u_{2,i},\ldots,u_{k,i})(p_{n}(u_{2,i},\ldots,u_{k,i}))^{n-1}=r_{n-\ell}(u_{2,i},\ldots,u_{k,i})(p_{n}(u_{2,i},\ldots,u_{k,i}))^{n-\ell}.$
Thus, for each $1\leq i\leq T$, $(u_{2,i},\ldots,u_{k,i})\in{\mathbb{F}}_{q}^{k-1}$ is a zero of the polynomial
$g_{\ell}(x_{2},\ldots,x_{k}):=p_{n-\ell}(x_{2},\ldots,x_{k})(p_{n}(x_{2},\ldots,x_{k}))^{n-1}-r_{n-\ell}(x_{2},\ldots,x_{k})(p_{n}(x_{2},\ldots,x_{k}))^{n-\ell}.$
Since $p_{n}$ and $p_{n-\ell}$ have degree at most $d$, and $r_{n-\ell}$ has degree at most $\ell d$, it follows that $g_{\ell}$ has degree at most $nd$. In view of the Schwartz-Zippel lemma, to conclude that $|Z|=T\leq ndq^{k-2}$, it suffices to prove the following claim.
Claim 3.4.
There exists $\ell_{0}$ with $1\leq\ell_{0}\leq n$ so that $g_{\ell_{0}}(x_{2},\ldots,x_{k})$ is not identically zero.
Proof of claim.
We give a proof by contradiction. Assume that $g_{\ell}(x_{2},\ldots,x_{k})$ is identically zero for any $\ell\in\{1,\ldots,n\}$. Then it follows from equation (3.5) that
$(p_{n}(x_{2},\ldots,x_{k}))^{n-1}f(x_{1},x_{2},\ldots,x_{k})=h^{2}(x_{1},x_{2},\ldots,x_{k}).$
We can write $f(x_{1},x_{2},\ldots,x_{k})=f_{1}(x_{1},x_{2},\ldots,x_{k})(f_{2}(x_{1},x_{2},\ldots,x_{k}))^{2}$ such that $f_{1}(x_{1},x_{2},\ldots,x_{k})$ is square-free. Note that $f_{1}$ still satisfies the primitive condition. Then we have
$(p_{n}(x_{2},\ldots,x_{k}))^{n-1}f_{1}(x_{1},x_{2},\ldots,x_{k})=(h(x_{1},x_{2},\ldots,x_{k})/f_{2}(x_{1},x_{2},\ldots,x_{k}))^{2}.$
Since $f_{1}$ is square-free, we must have $f_{1}(x_{1},x_{2},\ldots,x_{k})\mid(p_{n}(x_{2},\ldots,x_{k}))^{n-1}$. Thus,
$f_{1}(x_{1},x_{2},\ldots,x_{k})=g(x_{2},\ldots,x_{k})$
for some polynomial $g$. By the primitive condition, $g$ has no zero over $\overline{{\mathbb{F}}_{q}}$. This forces $g$ to be a constant polynomial and thus $f(x_{1},x_{2},\ldots,x_{k})=C(f_{2}(x_{1},x_{2},\ldots,x_{k}))^{2}$ for some $C\in{\mathbb{F}}_{q}$, violating our assumption. ∎
Remark 3.5.
Alternatively, the bound given in the statement of Proposition 3.1 can be rephrased in terms of $\ell,d$, and $n$. We note that $\deg p_{n}=d-n$, $\deg p_{n-\ell}=d-n+\ell$, and $\deg r_{n-\ell}\leq\ell d$. Hence, the degree of $g_{\ell}$ is bounded by
$\max\{(d-n+\ell)+(n-1)(d-n),\ell d+(n-\ell)(d-n)\}=n(d+\ell-n).$
The application of the Schwartz-Zippel lemma gives the bound $|Z|\leq n(d+\ell-n)q^{k-2}$. Combined with the bound on $|Y|\leq(d-n)q^{k-2}$, we have the following bound
$|X|=|Y|+|Z|\leq(nd+n\ell-n^{2}+d-n)q^{k-2}.$
In particular, when $n=d$ (e.g., when $f$ is in a diagonal form), we have the bound $|X|\leq n\ell q^{k-2}\leq d^{2}q^{k-2}$, which is better than the bound stated in Proposition 3.1.
Remark 3.6.
When $f(x_{1},x_{2},\ldots,x_{k})\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$ is diagonal and symmetric with even degree $d$, that is, $f(x_{1},x_{2},\ldots,x_{k})=ax_{1}^{d}+\cdots+ax_{k}^{d}$ with $a\in{\mathbb{F}}_{q}^{*}$, well-known estimates on the number of solutions of the diagonal equations can give bounds on $|X|$. In this case, we have
$X$ =$=\{(u_{2},\ldots,u_{k})\in{\mathbb{F}}_{q}^{k-1}~{}|~{}au_{2}^{d}+\cdots+au_{k}^{d}=0\}.$
For instance, [27, Theorem 6.36] gives $|X|=q^{k-2}+O_{d}(q^{k/2}).$ We can observe that the bound given in Proposition 3.1 is sharp up to the constant factor $d^{2}+d$.
4. Proof of Theorem 1.3
4.1. A sufficient condition
Now we provide the following sufficient condition on the quasirandomness of $Y_{f,q}$.
Proposition 4.1.
Let $f\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{2k}]$ be a symmetric polynomial of degree $d$. Let $\mathcal{B}$ be the set of tuples $(u_{2}(0),u_{2}(1),\ldots,u_{k}(0),u_{k}(1))\in{\mathbb{F}}_{q}^{2k-2}$ such that
$\prod_{\begin{subarray}{c}\varepsilon_{i}\in\{0,1\}\\ 2\leq i\leq k\end{subarray}}f\big{(}x,u_{2}(\varepsilon_{2}),u_{3}(\varepsilon_{3}),\ldots,u_{k}(\varepsilon_{k})\big{)}$
is a constant multiple of the square of a polynomial in ${\mathbb{F}}_{q}[x]$. Then
$N_{Y_{f,q}}^{*}(\mathcal{O}_{k}^{e})=\frac{q^{2k}}{2}+O_{d}(q^{2k-1}+|\mathcal{B}|q^{2}).$
Our proof relies on the celebrated Weil bound for complete character sums; see, for example, [27, Theorem 5.41].
Lemma 4.2 (Weil’s bound).
Let $\chi$ be a multiplicative character of ${\mathbb{F}}_{q}$ of order $e>1$, and let $g\in{\mathbb{F}}_{q}[x]$ be a monic polynomial of positive degree that is not an $e$-th power of a polynomial. Let $s$ be the number of distinct roots of $g$ in its splitting field over ${\mathbb{F}}_{q}$. Then for any $a\in{\mathbb{F}}_{q}$,
$\bigg{|}\sum_{x\in\mathbb{F}_{q}}\chi\big{(}ag(x)\big{)}\bigg{|}\leq(s-1)\sqrt{q}.$
Proof of Proposition 4.1.
Let $\chi$ be the quadratic character of ${\mathbb{F}}_{q}$. Let $v_{1},v_{2},\ldots,v_{k}$ be distinct vertices of $Y_{f,q}$. Then by definition, $\{v_{1},v_{2},\ldots,v_{k}\}$ forms an edge in $Y_{f,q}$ if and only if $\chi(f(v_{1},v_{2},\ldots,v_{k}))\in\{0,1\}$. Thus, $2k$ distinct vertices $u_{1}(0),u_{1}(1),\ldots,u_{k}(0),u_{k}(1)$ forms an induced copy of an EPO if and only if there are even number of vectors $\mathcal{E}=(\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{k})\in\{0,1\}^{k}$ such that
$\chi\big{(}f(u_{1}(\varepsilon_{1}),u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k}))\big{)}=-1.$
We then consider two cases:
Case 1: $f(u_{1}(\varepsilon_{1}),u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k}))=0$ for some $\mathcal{E}\in\{0,1\}^{k}$. By the Schwartz-Zippel lemma, the number of such induced copies of EPOs is at most $2^{k}dq^{2k-1}$.
Case 2: $f(u_{1}(\varepsilon_{1}),u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k}))\neq 0$ for all $\mathcal{E}\in\{0,1\}^{k}$. In this case,
$\prod_{\mathcal{E}\in\{0,1\}^{k}}\chi\big{(}f(u_{1}(\varepsilon_{1}),u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k})))\big{)}=(-1)^{r},$
where $r$ is the number of non-edges. Thus, we have
$\frac{1}{2}\bigg{(}1+\prod_{\mathcal{E}\in\{0,1\}^{k}}\chi\big{(}f(u_{1}(\varepsilon_{1}),u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k}))\big{)}\bigg{)}$
=$=\begin{cases}1,&\text{if }u_{1}(0),u_{1}(1),\ldots,u_{k}(0),u_{k}(1)\text{ forms an induced copy of an EPO}\\ 0,&\text{otherwise}.\end{cases}$
Therefore, we deduce that
$N_{Y_{f,q}}^{*}(\mathcal{O}_{k}^{e})=\sum\mathop{{}^{\prime}}\frac{1}{2}\bigg{(}1+\prod_{\mathcal{E}\in\{0,1\}^{k}}\chi\big{(}f(u_{1}(\varepsilon_{1}),u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k}))\big{)}\bigg{)}+O_{d}(q^{2k-1}),$
where the primed sum runs over all $2k$-tuples $(u_{1}(0),u_{1}(1),\ldots,u_{k}(0),u_{k}(1))\in{\mathbb{F}}_{q}^{2k}$ with distinct entries. Note that the number of $2k$-tuples $(u_{1}(0),u_{1}(1),\ldots,u_{k}(0),u_{k}(1))\in{\mathbb{F}}_{q}^{2k}$ with repeated entries is at most $\binom{2k}{2}q^{2k-1}$. Therefore, we still have
$N_{Y_{f,q}}^{*}(\mathcal{O}_{k}^{e})=\sum\frac{1}{2}\bigg{(}1+\prod_{\mathcal{E}\in\{0,1\}^{k}}\chi\big{(}f(u_{1}(\varepsilon_{1}),u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k}))\big{)}\bigg{)}+O_{d}(q^{2k-1}),$
where the sum runs over all $2k$-tuples $(u_{1}(0),u_{1}(1),\ldots,u_{k}(0),u_{k}(1))\in{\mathbb{F}}_{q}^{2k}$. It follows that $N_{Y_{f,q}}^{*}(\mathcal{O}_{k}^{e})=\frac{q^{2k}}{2}+S+O_{d}(q^{2k-1})$, where $S$ is the following sum below running over all $2k$-tuples $(u_{1}(0),u_{1}(1),\ldots,u_{k}(0),u_{k}(1))\in{\mathbb{F}}_{q}^{2k}$:
$S:=\sum\prod_{\mathcal{E}\in\{0,1\}^{k}}\chi\big{(}f(u_{1}(\varepsilon_{1}),u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k}))\big{)}.$
By changing the order of summation, we have
$S=$ $\sum_{\begin{subarray}{c}u_{2}(0),u_{2}(1)\\ ,\ldots,\\ u_{k}(0),u_{k}(1)\end{subarray}}\sum_{u_{1}(0)}\sum_{u_{1}(1)}\left(\prod_{\begin{subarray}{c}\varepsilon_{1}=0\\ \mathcal{E}\in\{0,1\}^{k}\end{subarray}}\chi\big{(}f(u_{1}(\varepsilon_{1}),u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k}))\big{)}\prod_{\begin{subarray}{c}\varepsilon_{1}=1\\ \mathcal{E}\in\{0,1\}^{k}\end{subarray}}\chi\big{(}f(u_{1}(\varepsilon_{1}),u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k}))\big{)}\right)$
$=$ $\sum_{\begin{subarray}{c}u_{2}(0),u_{2}(1)\\ ,\ldots,\\ u_{k}(0),u_{k}(1)\end{subarray}}\underbrace{\left(\sum_{u_{1}(0)}\chi\bigg{(}\prod_{\begin{subarray}{c}\varepsilon_{1}=0\\ \mathcal{E}\in\{0,1\}^{k}\end{subarray}}f\big{(}u_{1}(0),u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k})\big{)}\bigg{)}\right)}_{(1)}\underbrace{\left(\sum_{u_{1}(1)}\chi\bigg{(}\prod_{\begin{subarray}{c}\varepsilon_{1}=1\\ \mathcal{E}\in\{0,1\}^{k}\end{subarray}}f\big{(}u_{1}(1),u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k})\big{)}\bigg{)}\right)}_{(2)}.$
Note that, if we fix $(u_{2}(0),u_{2}(1),\ldots,u_{k}(0),u_{k}(1))\in{\mathbb{F}}_{q}^{2k-2}$, then
$g(x):=\prod_{\begin{subarray}{c}\varepsilon_{1}=0\\ \mathcal{E}\in\{0,1\}^{k}\end{subarray}}f(x,u_{2}(\varepsilon_{2}),\ldots,u_{k}(\varepsilon_{k}))$
is a polynomial in $x$ of degree at most $2^{k-1}d$ ( also when $\epsilon_{1}=1$). Thus, the absolute value of both $(1)$ and $(2)$ is at most $q$ provided that $g(x)$ is a constant multiple of the square of a polynomial in ${\mathbb{F}}_{q}[x]$; otherwise, Weil’s bound implies that the absolute value of both $(1)$ and $(2)$ is at most $2^{k-1}d\sqrt{q}$. Thus, we can bound $|S|$ by splitting the sum over $\mathcal{B}$ and ${\mathbb{F}}_{q}^{2k-2}\setminus\mathcal{B}$ to conclude that
$|S|\leq|\mathcal{B}|q^{2}+(q^{2k-2}-|\mathcal{B}|)(2^{k-1}d\sqrt{q})^{2}\leq|\mathcal{B}|q^{2}+q^{2k-2}(2^{k-1}d\sqrt{q})^{2}=O_{d}(q^{2k-1}+|\mathcal{B}|q^{2}),$
as required. ∎
4.2. Proof of Theorem 1.3 via Proposition 4.1
Let $Y_{f,q}\in\mathcal{X}_{d}$; we claim that
$N^{*}_{Y_{f,q}}(\mathcal{O}_{k}^{e})=\frac{q^{2k}}{2}+O_{d}(q^{2k-1}).$
As in Proposition 4.1, let $\mathcal{B}$ be the set of tuples $(u_{2}(0),u_{2}(1),\ldots,u_{k}(0),u_{k}(1))\in{\mathbb{F}}_{q}^{2k-2}$ such that
$\prod_{\begin{subarray}{c}\varepsilon_{i}\in\{0,1\}\\ 2\leq i\leq k\end{subarray}}f\big{(}x,u_{2}(\varepsilon_{2}),u_{3}(\varepsilon_{3}),\ldots,u_{k}(\varepsilon_{k})\big{)}$ (4.1)
is a constant multiple of the square of a polynomial in ${\mathbb{F}}_{q}[x]$. In view of Proposition 4.1, to complete the proof of Theorem 1.3, it suffices to show that $|\mathcal{B}|=O_{d}(q^{2k-3}).$ Since $f$ satisfies the primitive condition, by definition, the polynomial $f(x,v_{2},v_{3},\ldots,v_{k})$ is not identically zero for any $(v_{2},v_{3},\ldots,v_{k})\in\overline{{\mathbb{F}}_{q}}^{k-1}$. This fact will be used repeatedly throughout the proof below. In particular, this fact implies that the polynomial in 4.1 is not identically zero for any $(u_{2}(0),u_{2}(1),\ldots,u_{k}(0),u_{k}(1))\in{\mathbb{F}}_{q}^{2k-2}$.
We first give an upper bound on the number of tuples $(u_{2}(0),u_{2}(1),\ldots,u_{k}(0),u_{k}(1))\in\mathcal{B}$ such that $f(x,u_{2}(0),u_{3}(0),\ldots,u_{k}(0))$ is a constant multiple of the square of a polynomial in ${\mathbb{F}}_{q}[x]$. By Proposition 3.1, the number of such tuples is at most $O_{d}(q^{k-2})$, which is hence upper-bounded by $O_{d}(q^{2k-3})$. It remains to bound the number of tuples $(u_{2}(0),u_{2}(1),\ldots,u_{k}(0),u_{k}(1))$ in $\mathcal{B}$ such that $f(x,u_{2}(0),u_{3}(0),\ldots,u_{k}(0))$ is not a constant multiple of the square of a polynomial in ${\mathbb{F}}_{q}[x]$. Note that for such a tuple in $\mathcal{B}$, since the polynomial in 4.1 is not identically zero, there exist $\varepsilon_{2},\ldots,\varepsilon_{k}\in\{0,1\}$ that are not all zero, such that
$\gcd\big{(}f(x,u_{2}(0),\ldots,u_{k}(0)),f(x,u_{2}(\varepsilon_{2}),u_{3}(\varepsilon_{3}),\ldots,u_{k}(\varepsilon_{k}))\big{)}$
has degree at least $1$; equivalently, these two polynomials have a common root $c\in\overline{{\mathbb{F}}_{q}}$. Note that for each $(u_{2}(0),u_{3}(0),\ldots,u_{k}(0))\in{\mathbb{F}}_{q}^{k-1}$, the polynomial $f(x,u_{2}(0),\ldots,u_{k}(0))$ has at most $d$ roots in $\overline{{\mathbb{F}}_{q}}$ since it is not identically zero. Thus, it suffices to show that if $u_{2}(0),\ldots,u_{k}(0)\in{\mathbb{F}}_{q}$ are fixed, and $c\in\overline{{\mathbb{F}}_{q}}$ is fixed, then for each $\varepsilon_{2},\ldots,\varepsilon_{k}\in\{0,1\}$ not all zero, the number of tuples $(u_{2}(1),u_{3}(1),\ldots,u_{k}(1))\in{\mathbb{F}}_{q}^{k-1}$ such that $f(c,u_{2}(\varepsilon_{2}),u_{3}(\varepsilon_{3}),\ldots,u_{k}(\varepsilon_{k}))=0$ is $O_{d}(q^{k-2})$.
Let $u_{2}(0),u_{3}(0),\ldots,u_{k}(0)\in{\mathbb{F}}_{q}$, and $c\in\overline{{\mathbb{F}}_{q}}$ be all fixed. Let $\varepsilon_{2},\ldots,\varepsilon_{k}\in\{0,1\}$ such that they are not all zero. Since $f$ is symmetric, without loss of generality, we can assume that there is $t\geq 2$, such that $\varepsilon_{2}=\varepsilon_{3}=\cdots=\varepsilon_{t}=1$ and $\varepsilon_{t+1}=\varepsilon_{t+2}=\cdots=\varepsilon_{k}=0$. Then
$f(x_{2},x_{3},\ldots,x_{t},c,u_{t+1}(0),u_{t+2}(0),\ldots,u_{k}(0))=f(c,x_{2},x_{3},\ldots,x_{t},u_{t+1}(0),u_{t+2}(0),\ldots,u_{k}(0))$
is a polynomial with $t-1$ variables defined over $\overline{{\mathbb{F}}_{q}}$, and it is not identically zero since $t\geq 2$. Thus, by the Schwarz-Zippel lemma, the number of tuples $(u_{2}(1),u_{3}(1),\ldots,u_{t}(1))\in{\mathbb{F}}_{q}^{t-1}$ such that $f(c,u_{2}(1),u_{3}(1),\ldots,u_{t}(1),u_{t+1}(0),u_{t+2}(0),\ldots,u_{k}(0))=0$ is at most $dq^{t-2}$. Since $u_{t+1}(1),u_{t+2}(1),\ldots,u_{k}(1)$ can be chosen arbitrarily in ${\mathbb{F}}_{q}$, it follows that the number of tuples $(u_{2}(1),u_{3}(1),\ldots,u_{k}(1))\in{\mathbb{F}}_{q}^{k-1}$ such that
$f(c,u_{2}(\varepsilon_{2}),u_{3}(\varepsilon_{3}),\ldots,u_{k}(\varepsilon_{k}))=f(c,u_{2}(1),u_{3}(1),\ldots,u_{t}(1),u_{t+1}(0),u_{t+2}(0),\ldots,u_{k}(0))=0$
is at most $dq^{t-2}\cdot q^{k-t}=dq^{k-2}=O_{d}(q^{k-2})$, as required.
Remark 4.3.
Theorem 1.3 in fact applies to a slightly larger family of polynomials. Let $f$ be a symmetric polynomial of degree $d$ in ${\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$ that is not necessarily admissible. Assume that $f$ can be written as $g\cdot h$, where $g$ is a symmetric polynomial that can be written as a product of polynomials depending at most $k-1$ variables, and $h$ is an admissible polynomial. For such an $f$, one can begin with the product in 4.1 and apply a similar argument to show the same result for $N^{*}_{Y_{f,q}}(\mathcal{O}_{k}^{e})$.
5. Asymptotic formula for the number of $f$-Diophantine sets with size $m$
Recently, Slavov [31, Theorem 3 and Remark 5] used an explicit version of the Lang-Weil bound [6, Theorem 7.1] to prove a result on the joint distribution of square values of polynomials over a finite field. The following lemma is a special case of his result.
Lemma 5.1 (Slavov).
Let $n,m,d$ be positive integers and $q$ be an odd prime power. Let $f_{1},f_{2},\ldots,f_{n}\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{m}]$ with degree $d$. Then the following are equivalent:
(i) For any nonempty subset $T\subset\{1,2,\ldots,n\}$, the polynomial $\prod_{i\in T}f_{i}$ is not a constant multiple of the square of a polynomial in ${\mathbb{F}}_{q}[x_{1},\ldots,x_{m}]$.
(ii) The number of $(a_{1},a_{2},\ldots,a_{m})\in{\mathbb{F}}_{q}^{m}$ such that $f_{i}(a_{1},a_{2},\ldots,a_{m})$ is a nonzero square in ${\mathbb{F}}_{q}$ for all $1\leq i\leq n$ is
$\frac{q^{m}}{2^{n}}+O(d^{2n}q^{m-1/2}+(2d)^{13n/3}q^{m-1}),$
where the implied constant in the error term is absolute.
Note that condition (i) stated in the above lemma is slightly different from the condition (i) stated originally in [31, Theorem 3], but clearly they are equivalent.
We end the paper by presenting a proof of Theorem 1.7. Most arguments used in the proof below have already appeared in previous sections.
Proof of Theorem 1.7.
Let $q$ be an odd prime power and $f\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{k}]$ be an admissible polynomial with degree $d$. For each $m\geq k$, let $N(f,m)$ be the number of $f$-Diophantine sets of size $m$.
Let $\chi$ be the quadratic character of ${\mathbb{F}}_{q}$, and let $\tilde{\chi}:{\mathbb{F}}_{q}\to{\mathbb{F}}_{q}$ such that $\tilde{\chi}(x)=\chi(x)$ for $x\neq 0$, and $\tilde{\chi}(0)=1$. Similar to the proof of Proposition 4.1, we can express $N(f,m)$ in terms of the following character sum:
$N(f,m)=\sum_{\begin{subarray}{c}A\subset\mathbb{F}_{q}\\ |A|=m\end{subarray}}\prod_{\begin{subarray}{c}T=\{b_{1},\ldots,b_{k}\}\subset A\\ |T|=k\end{subarray}}\bigg{(}\frac{1+\tilde{\chi}\big{(}f(b_{1},\ldots,b_{k})\big{)}}{2}\bigg{)}=m!\cdot S(f,m)+O(dq^{m-1}),$
where
$S(f,m):=\sum\mathop{{}^{\prime}}\prod_{\begin{subarray}{c}I=\{i_{1},i_{2},\ldots,i_{k}\}\subset\{1,2,\ldots,m\}\\ |I|=k\end{subarray}}\bigg{(}\frac{1+\chi\big{(}f(a_{i_{1}},\ldots,a_{i_{k}})\big{)}}{2}\bigg{)},$ (5.1)
with the primed sum runs over all $(a_{1},a_{2},\ldots,a_{m})\in{\mathbb{F}}_{q}^{m}$ such that $f(a_{i_{1}},\ldots,a_{i_{k}})\neq 0$ for all $I=\{i_{1},i_{2},\ldots,i_{k}\}\subset\{1,2,\ldots,m\}$ with $|I|=k$. It follows that
$N(f,m)=\frac{S(f,m)}{m!}+O(dq^{m-1}).$ (5.2)
Let $\mathcal{I}$ be the collection of all subsets of $\{1,2,\ldots,m\}$ of size $k$. For each $I=\{i_{1},i_{2},\ldots,i_{k}\}\in\mathcal{I}$, define $f_{I}\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{m}]$ by $f_{I}(x_{1},x_{2},\ldots,x_{m})=f(x_{i_{1}},x_{i_{2}},\ldots,x_{i_{k}})$. Based on equation (5.1), we see that $S(f,m)$ precisely counts the number of $(a_{1},a_{2},\ldots,a_{m})\in{\mathbb{F}}_{q}^{m}$ such that $f_{I}(a_{1},a_{2},\ldots,a_{m})$ is a nonzero square in ${\mathbb{F}}_{q}$ for all $I\in\mathcal{I}$.
Next, we show that for each nonempty subset $T\subset\mathcal{I}$, the polynomial $\prod_{I\in T}f_{I}$ is not a constant multiple of the square of a polynomial in ${\mathbb{F}}_{q}[x_{1},\ldots,x_{m}]$. When $|T|=1$, this is obvious from the definition of admissible polynomials. Suppose otherwise that there is a subset $T\subset\mathcal{I}$ with $|T|\geq 2$ such that the polynomial $\prod_{I\in T}f_{I}$ is a constant multiple of the square of a polynomial in ${\mathbb{F}}_{q}[x_{1},\ldots,x_{m}]$. Pick $J\in T$. From the definition of admissible polynomials, $f_{J}$ is not a constant multiple of the square of a polynomial in ${\mathbb{F}}_{q}[x_{1},\ldots,x_{m}]$. Thus, there is $J^{\prime}\in T$ with $J^{\prime}\neq J$, such that $\gcd(f_{J},f_{J^{\prime}})$ is a polynomial with degree at least $1$. It follows that $J\cap J^{\prime}\neq\emptyset$. Without loss of generality, assume that $J\cap J^{\prime}=\{1,2,\ldots,t\}$ with $11\leq t\leq k-$ 1. Then $\gcd(f_{J},f_{J^{\prime}})=g(x_{1},x_{2},\ldots,x_{t})$ for some polynomial $g\in{\mathbb{F}}_{q}[x_{1},x_{2},\ldots,x_{t}]$ with degree $1$, and in particular $g$ is a factor of $f$. However, this violates the primitive condition of the definition of admissible polynomials.
Thus, by setting $n=\binom{m}{k}$ applied to the collection of polynomials $\{f_{I}:I\in\mathcal{I}\}$, we have verified condition (i) in Lemma 5.1, and thus (ii) in the lemma implies that
$S(f,m)=\frac{q^{m}}{2^{\binom{m}{k}}}+O\big{(}d^{2\binom{m}{k}}q^{m-1/2}+(2d)^{13\binom{m}{k}/3}q^{m-1}\big{)}.$ (5.3)
Combining equations (5.2) and (5.3), we obtain the desired result. ∎
Acknowledgments
The authors thank Jaehoon Kim and Joonkyung Lee for their valuable feedback on a preliminary version of the manuscript. The third author thanks Hyunwoo Lee for helpful discussions on the quasi-randomness of hypergraphs. The research of the second author was supported in part by an NSERC fellowship. The third author was supported by the Institute for Basic Science (IBS-R029-C1).